OpenStudy (anonymous):

4x^(3)-8x^(2)+2x-1 over 2x

OpenStudy (anonymous):

\[\large \frac{4x^3-8x^2+2x-1}{2x}=\frac{4x^3}{2x}-\frac{8x^2}{2x}+\frac{2x}{2x}-\frac{1}{2x} \] can you simplify from here?

OpenStudy (anonymous):

 hope that helps :)